# Interaction-Aware Model Predictive Decision-Making for Socially-Compliant Autonomous Driving in Mixed Urban Traffic Scenarios

### v1

@@ 6 Summary
This paper proposes an interaction-aware model predictive decision-making system and its real-time implementation for automated vehicles interacting with pedestrians in urban, low-speed scenarios. Furthermore, we designed a human-in-the-decision-loop study to examine the human-automation action-reaction cycle. The human-in-the-decision-loop study, conducted with 25 participants, demonstrated that intention-aware decision-making algorithms lead to faster resolution of human-automation negotiations and improved subjective evaluations. These findings highlight the significance of intention-aware decision-making for autonomous vehicles in urban settings. Our future work will feature an adaptive decision-making algorithm based on Varga (2024). Furthermore, we plan to integrate a probabilistic estimation of human crossing intentions by modeling their stochastic behavior, see Škugor et al. (2023).

### v2

@@ 6 Summary
This paper proposes an interaction-aware model predictive decision-making system and its real-time implementation for automated vehicles interacting with pedestrians in urban, low-speed scenarios. Furthermore, we designed a human-in-the-decision-loop study to examine the human-automation action-reaction cycle. The human-in-the-decision-loop study,conducted with 25 participants, demonstrated that intention-aware decision-making algorithms lead to faster resolution of human-automation negotiations and improved subjective evaluations. These findings highlight the significance of intention-aware decision-making for autonomous vehicles in urban settings. Our future work will feature an adaptive decision-making algorithm based on Varga [2024]. Furthermore, we plan to integrate a probabilistic estimation of human crossing intentions by modeling their stochastic behavior, see Škugor et al. [2023].
@@
Declaration of generative AI and AI-assisted technologies in the manuscript preparation process
During the preparation of this work, the authors used ChatGPT (OpenAI) to assist with improving language, readability, and organization of parts of the manuscript. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.